A 13ft. ladder is leaning against a building. Thhe bottom of the ladder is 5ft from the building. How high is the top of the ladder?
Answer by checkley75(3666)

5^2+x^2=13^2
25+x^2=169
x^2=169-25
x^2=144
x=sqrt144
x=12 height of the ladder against the building.